OpenStudy (anonymous):

Solve the open sentence. –3 ≤ n + 6 ≤ 7 (Points : 1) n ≤ –9 and n ≤ 6 n ≥ –9 and n ≤ 1 n ≤ 3 and n ≤ 7 n ≥ 9 and n ≤ 7

OpenStudy (anonymous):

First subtract 6 from both sides of the operator \( -3 \leq n + 6 \leq 7 \) \( -3-6 \leq n + 6 -6 \leq 7 -6\) solve that and show me what you have.